OpenStudy (lukecrayonz):

Using DeMoivre's Theorem, find the indicated power of the complex number. Express the result in standard form. (2+2i)^6

OpenStudy (amistre64):

Demovers is what: r^n (cos(nax) + sin(nbx)) ?? im sure i never remember that one right

OpenStudy (amistre64):

the theta part is what i was missing :/

OpenStudy (amistre64):

given:\[(a+bi)^n\] \[r=\sqrt{a^2+b^2};\ t=tan^{-1}(\frac{b}{a})\] \[(a+bi)^n=r^n(cos(nt)+isin(nt))\]

OpenStudy (lukecrayonz):

Amistre

OpenStudy (lukecrayonz):

@amistre64

OpenStudy (lukecrayonz):

Need your help real quick

OpenStudy (amistre64):

timmys in the well?

OpenStudy (lukecrayonz):

So (2+2i)^6, sqrt(8)=2sqrt(2), and tan^-1=2/2 = pi/4, now what?

OpenStudy (lukecrayonz):

Lassie!

OpenStudy (amistre64):

now we fill in demovers

OpenStudy (lukecrayonz):

so 2sqrt(2)^6(cos(5*pi/4)+isin(5*pi/4))?

OpenStudy (amistre64):

\[(a+bi)^n=r^n(cos(nt)+isin(nt))\] \[(2+2i)^6=(2\sqrt{2})^6(cos(6\frac{pi}4)+isin(6\frac{pi}4))\]

OpenStudy (amistre64):

\[(2\sqrt{2})^6(cos(6\frac{pi}4)+isin(6\frac{pi}4))\] \[(64*8)(cos(\frac{3pi}2)+isin(\frac{3pi}2 ))\]

OpenStudy (lukecrayonz):

Sorry I meant 6 in my equation, typoed.

OpenStudy (amistre64):

sin(t) = -1 there and cos(t)=0

OpenStudy (amistre64):

(64*8)(-i)

OpenStudy (amistre64):

480 32 ---- 512 -512i

OpenStudy (lukecrayonz):

Is there a different more easier way to do this?

OpenStudy (lukecrayonz):

and what does "sin(t) = -1 there and cos(t)=0" have to do with anything?